OpenStudy (anonymous):

Use the Five steps. The premiter of a certain rectangle is no more than 94 feet. If the length is eleven feet more than width, what is the width?

ganeshie8 (ganeshie8):

2(l+w) <= 94 2(w+11+w) <=94 2(2w+11) < = 94 4w <= 94-22 = 72 w <= 18

OpenStudy (anonymous):

thank you

ganeshie8 (ganeshie8):

yw..